OpenStudy (anonymous):

what is the least number that can be exactly divided by 56, 63 and 84?

OpenStudy (curry):

she is asking for the lcm i bleive not lcf

OpenStudy (anonymous):

curry is right.

OpenStudy (zarkon):

504

OpenStudy (curry):

zarkon help me

OpenStudy (anonymous):

To find the lcm of two numbers a and b: Simply divide (a * b) by their gcd

OpenStudy (anonymous):

slip on my part.

OpenStudy (zarkon):

lcm(a,b,c)=lcm(a,lcm(b,c))

OpenStudy (anonymous):

for all those numbers up their, their gcd is 7. lcm(56, 63) = 504 lcm(84, 63) = 252 it becomes clear that lcm(504, 252) is 504

OpenStudy (anonymous):

there* I can't believe I misspelled that :(

OpenStudy (asnaseer):

another way to look at this is to write each number as a product of their prime factors, so: \[56=2^3*7\]\[63=3^2*7\]\[84=2^2*3*7\] then just select every unique prime factor (picking the one with the largest power if two or more numbers share a prime) and multiply these together to get your Lowest Common Multiple - see diagram:  so LCM(56, 63, 84) is:\[2^3*3^2*7=504\]